A rectangular flower garden in a park is 30 feet wide and 40 feet long. A sidewalk around the perimeter of the garden is being planned. The gardener has enough money to pour 624 feet of cement sidewalk. Find the width of the sidewalk.
Answer by friesr(113)

The parimeter of the garden is 140ft.. Assuming you actually ment 624ft^2 of sidewalk then 624/140 = 156/35 or approx 4.45ft.
---
140 * (156/35) = 624